Mr.Ali has an income of Rs.30,000 and he purchases 100 units of X-GOOD. His income decreases by 20% and now he can purchase 90 units of X-GOOD.Calculate Calculate elasticity and also mention the type of elasticity.
1
Expert's answer
2020-04-30T09:59:04-0400

The income elasticity is:

"Ei = \frac{90 - 100}{24,000 - 30,000}\times\frac{24,000 + 30,000}{90 + 100} = 9/19," so X-GOOD is a normal good.